# Arecibo Wow! I: An Astrophysical Explanation for the Wow! Signal

### v1

@@ Abstract
The Ohio State University Big Ear radio telescope detected in 1977 the Wow! Signal, one of the most famous and intriguing signals of extraterrestrial origin. The signal was characterized by its strong relative intensity and narrow bandwidth near the 1420 MHz hydrogen line, but its source has never been detected again despite numerous follow-up attempts. Arecibo Wow! is a new project that aims to find similar signals in archived data from the Arecibo Observatory. From 2017 to 2020, we observed many targets of interest at 1 to 10 GHz with the 305-meter telescope. We continued these observations at 8 GHz with the 12-meter telescope in 2023. Most of these targets were red dwarf stars with planets, some potentially habitable. Here we present our first results of drift scans made between February and May 2020 at 1420 MHz. The methods, frequency, and bandwidth of these observations are similar to those used to detect the Wow! Signal. However, our observations are more sensitive, have better temporal resolution, and include polarization measurements. We report the detection of narrowband signals ($\Delta\nu\leq$ 10 kHz) near the hydrogen line similar to the Wow! Signal, although two-orders of magnitude less intense and in multiple locations. Despite the similarities, these signals are easily identifiable as due to interstellar clouds of cold hydrogen (HI) in the galaxy. We hypothesize that the Wow! Signal was caused by sudden brightening from stimulated emission of the hydrogen line due to a strong transient radiation source, such as a magnetar flare or a soft gamma repeater (SGR). These are very rare events that depend on special conditions and alignments, where these clouds might become much brighter for seconds to minutes. The original source or the cloud might not be detectable, depending on the sensitivity of the telescope or because the maximum brightness might arrive seconds later to the observer. The precise location of the Wow! Signal might be determined by searching for transient radio sources behind the cold hydrogen clouds in the corresponding region. Our hypothesis explains all observed properties of the Wow! Signal, proposes a new source of false positives in technosignature searches, and suggests that the Wow! Signal could be the first recorded event of an astronomical maser flare in the hydrogen line.
Technosignatures — Astrobiology — ISM: clouds — Stars: Magnetars — Interdisciplinary Astronomy — History of Astronomy
facilities: Arecibosoftware: The Arecibo Software Library (https://www.naic.edu/arecibo/), IDLAstro (https://github.com/wlandsman/IDLAstro)
@@ 1 Introduction
The Wow! Signal, detected on August 15, 1977, by the Ohio State University’s Big Ear radio telescope, remains one of the most compelling potential instances of extraterrestrial communication (see e.g. Wright 2021 and references therein). The signal, identified by astronomer Jerry R. Ehman, exhibited a narrowband radio frequency of approximately 1420.456 MHz, which is notably close to the hydrogen line frequency, a region of the radio spectrum often considered a natural candidate for interstellar communication (Kraus et al., 1979). The signal lasted for 72 seconds, corresponding to the time it took the telescope beam to sweep past the source, but no follow-up observations have successfully replicated this anomaly (Gray & Ellingsen, 2002). The transient nature of the Wow! Signal, combined with its strength and frequency, has led to extensive debate and investigation within the SETI community (Search for Extraterrestrial Intelligence), yet its exact origin remains undetermined (see, e.g., Tarter 2001).
@@
Despite numerous efforts using various radio telescopes, including the Very Large Array (VLA), no repeat detections have been made at the original frequency or any other candidate frequencies (Backus et al., 2004; Perez et al., 2022). The search for similar signals continues to be a significant focus within the SETI community, employing advanced digital signal processing techniques and expanded observational capabilities to improve the chances of detecting faint and transient signals (Werthimer et al., 2001). The Allen Telescope Array (ATA) was also used to search for possible repetitions (Harp et al., 2020). The researchers monitored the specific frequency and sky coordinates associated with the original signal, using the ATA’s capabilities to capture a broad frequency range in search of potential extraterrestrial signals. Despite comprehensive observation efforts spanning multiple sessions, the study did not result in the detection of any repeat signals. This outcome suggests that the Wow! Signal may have been a unique occurrence or an artifact of terrestrial interference, rather than a signal of extraterrestrial origin.
The implications of these findings are significant for SETI research. They highlight the need for continuous long-term monitoring strategies and developing more sensitive detection technologies. The study also stresses the importance of revisiting and re-analyzing historical datasets with newer, more advanced technology, which could potentially uncover subtleties missed in earlier analyses. This approach could lead to a better understanding of the transient phenomena observed in radio astronomy and their origins. The absence of repetitions underscores the challenges in SETI, emphasizing the sporadic nature of potential non-terrestrial signals.
The main objective of this paper is to present the first results of our Arecibo Wow! Project (Sections 3 to 7). These results suggest an astrophysical explanation to the Wow! Signal that can be tested with follow-up observations. This paper also presents a comparison between the Big Ear and the Arecibo Telescope, which should provide a context for the detection of the Wow! Signal and our observations. In Section 2 we discuss the detection and interpretation of the Wow! Signal. In Section 3 we provide a brief history of SETI projects at the Arecibo Observatory. We present concluding remarks in Section 8.
2 The Wow! Signal
@@
In 1973, following a shift in funding priorities by the United States Congress, the National Science Foundation (NSF) withdrew its financial support for the Ohio Sky Survey, which had a significant impact on the project. The loss of funding led to the cessation of the Ohio Sky Survey and the eventual disbandment of the survey team. Despite these setbacks, the Big Ear radio telescope was repurposed for a new mission: the systematic search for narrow-band radio signals. Unlike wide-band radio sources, which are typically natural and generate emissions across a broad spectrum (including radio, optical, X-ray, and gamma-ray bands), narrow-band sources are often artificial, such as those used in AM, FM, TV broadcasts, satellite communications, and radar, implying potential intelligent origin. With the narrow-band search underway, a systematic survey of the declination range from +64° to -36° began (Dixon & Cole, 1977).
@@
The data collected by the Big Ear radio telescope were processed and recorded in a series of computer printouts, each row encapsulating approximately 12 seconds of sidereal time (see Figure 1). During each 10-second period, the telescope collected one intensity value per second across 50 channels, averaging these ten values to represent the intensity of each channel throughout the 10-second integration period (Dixon & Cole, 1977). The remaining 2 seconds of the 12-second interval were dedicated to computer analysis, where it processed the data and identified any notable phenomena. The printouts displayed the average intensities for each of the 50 channels (each having $\Delta nu=10$ kHz), arranged sequentially from left to right, with channel 1 on the left side and channel 50 on the right side.
Figure 1: Full page of the computer printout with the Wow! Signal labeled in red. Credit: The Ohio History Connection Collections.
@@
Intensity values were converted to a single character for display: zero values were rendered as blanks, integers 1 to 9 were printed directly as digits, and integers 10 to 35 were represented by uppercase letters A through Z, respectively. Each character corresponds to a specific range of signal-to-noise ratios: “6” represents a ratio between 6 and 7, “E” corresponds to a ratio between 14 and 15, “Q” to 26 and 27, “U” to 30 and 31, “J” to 19 and 20, and “5” to 5 and 6. During a standard analysis of the data gathered on August 15, an unusual character string, “6EQUJ5,” was detected in channel 2 of the printout on August 19, 1977. This sequence was immediately recognized as indicative of a narrow-band radio source with a small angular diameter, a potential signal of intelligent origin. The strongest signal, represented by the character ”U,” indicates that the signal was 30.5 $\pm$0.5 times stronger than the background noise. This signal was later referred to as the Wow! Signal. In Table 1 we describe the characteristics of the Wow! Signal, including location, time, and absolute intensity.
@@
Frequency 1420.4556 MHz
Bandwidth narrow-band ($\leq$ 10 kHz)
Signal Strength 30 times background noise (SNR $\approx$30)
@@
Positive Horn Coordinates (Equatorial) RA: 19h 25m 31s ± 10s, Dec: -26d 57m ± 20m (J2000)
Negative Horn Coordinates (Equatorial) RA: 19h 28m 22s ± 10s, Dec: -26d 57m ± 20m (J2000)
Positive Horn Coordinates (Galactic) lon: 11.664d, lat: -18.916d
Negative Horn Coordinates (Galactic) lon: 11.909d, lat: -19.508d
@@
The hypothesis that has gained the most attention, especially among the public, is that the Wow! Signal was a transmission from an extraterrestrial civilization (Wright, 2021). The frequency of the signal, close to the hydrogen line, is considered a logical choice for interstellar communication because it is naturally quiet and would be universally recognized by technologically advanced civilizations (Drake & Sagan, 1973). However, this hypothesis faces critical issues. The signal was never repeated, despite numerous follow-up observations. In the context of SETI, repetition is essential for verification, as a one-time detection makes it difficult to rule out other sources. Moreover, the absence of modulation or encoding in the signal, which would be expected from an intentional transmission, further complicates this interpretation.
@@
Another possibility is that the Wow! Signal originated from a natural astrophysical source. Candidates include stellar flares, pulsars, or other cosmic phenomena that emit radio waves (for a recent explanation overview, see Kipping & Gray 2022). However, these sources typically exhibit characteristics such as periodicity or broader bandwidths, which the Wow! Signal lacks. The narrow-band nature of the signal is particularly unusual because natural emissions are generally broadband. Additionally, the hydrogen line frequency is not generally associated with natural radio emissions from known astrophysical objects, making this hypothesis less likely.
@@
Terrestrial interference is another plausible explanation. The signal could have been caused by a reflected terrestrial transmission or an Earth-based signal that temporarily overwhelmed the receiver. The 1420 MHz band is protected for astronomical purposes, but illegal or accidental transmission in this band could have occurred. However, the Big Ear radio telescope, which detected the Wow! Signal, had a design that helped filter out terrestrial sources, and the signal appeared only in one of the two feed horns, complicating the explanation of interference [^1] For a description and some of the hypotheses that have been raised about the terrestrial and extraterrestrial origin of the signal see the detailed account written by Jerry R. Ehman himself in 2010 that is still available at this url (last visited: August 16, 2024). No known terrestrial source has been identified that matches the characteristics of the Wow! Signal. Moreover, a reflecting piece of space debris must have been moving very slowly and not tumbling to mimic a point source in the celestial sphere, which is unlikely (Ehman, 2011a).
@@
One hypothesis suggests that the signal could have been caused by a reflection of Earth-based signals off space debris or satellites (Ehman, 2011b). This idea posits that a radio signal from Earth might have bounced off a satellite or space debris and was then picked up by the telescope. However, this explanation struggles with the same issue of non-repetition. The duration of the signal and its specific characteristics do not align neatly with the known satellite reflections or the behavior of the debris. Additionally, the Doppler shift expected from such a reflection was not observed.
@@
There is also the consideration of interstellar propagation effects, where the signal could have been distorted or focused through gravitational lensing or other space phenomena. These effects could potentially create a signal that mimics the characteristics of the Wow! Signal. However, the gravitational lensing hypothesis suffers from the drawback of the short signal duration – the Wow! Signal was not detected in the second horn of the Big Ear whose beam swept the same area of the sky only 5 minutes later (Ehman, 2011a). Additionally, such a scenario would likely involve complex interactions with multiple cosmic objects, which are improbable and would likely have been observed again under similar conditions.
@@
Finally, human error has been considered in the recording or processing of the data. Data handling methods in 1977 were not as sophisticated as today, and it is conceivable that a mistake or anomaly in the data processing chain produced the Wow! Signal. However, this explanation lacks specific evidence and does not account for the signal’s characteristics, which seem consistent with a real astronomical event.
@@
Table 2: General Comparison of the telescopes of the Arecibo Observatory (AO) and the Ohio State University Observatory (OSUO) in the L-band
@@
Location Arecibo, Puerto Rico Delaware, Ohio, USA
@@
Elevation 497 meters 276 meters
Type Spherical Reflector Cylindrical Paraboloid
Diameter 305 meters 110 meters
Illuminated Area 213 x 237 m ($\sim$73,000 m²) $\sim$8,000 m²
Radius of Curvature of Primary Reflector 265 m N/A
Operational Years 1963 - 2020 1963 - 1998
Maximum Declination Coverage -1 to +38 degrees +39 to +41 degrees
Data Collection Methods Fixed dish with moving receiver Fixed dish with Earth’s rotation
@@
Zenith Angle Range 0 to 19.7 degrees Fixed
Pointing Accuracy $\sim$1 arcsecond N/A
Table 3: Comparison of L-band receiver of the Arecibo Telescope (AO) and the Ohio State University Telescope (Big Ear).
Property Arecibo Telescope Big Ear
@@
Frequency Range 1.12 to 1.73 GHz 1.42 GHz (primary)
Frequency Channels 7 x 8192 50
Gain $\sim$10.5 K/J $\sim$0.1 K/J
System Equivalent Flux Density (SEFD) $\sim$2.4 Jy $\sim$50 Jy
Half-Power Beam Width (HPBW) (Az x ZA in) 3.1 x 3.5 arcminutes 8 x 40 arcminutes
Transit Time $\sim$48 seconds $\sim$180 seconds
3 SETI efforts from Arecibo

### v2

@@ Abstract
The Ohio State University Big Ear radio telescope detected in 1977 the Wow! Signal, one of the most famous and intriguing signals of extraterrestrial origin. Characterized by its strong relative intensity and narrow bandwidth near the 1420 MHz hydrogen line, its source has never been detected again despite numerous follow-up attempts. Arecibo Wow! is a new technosignature project using archived data from the Arecibo Observatory. Here we present our first results of drift scans made between February and May 2020 at 1420 MHz. The methods, frequency, and bandwidth of these observations are similar to those used to detect the Wow! Signal. However, our observations are more sensitive, have better temporal resolution, and include polarization measurements. We report the detection of narrowband signals ($\Delta\nu\leq$ 10 kHz) near the hydrogen line similar to the Wow! Signal, although two-orders of magnitude less intense and in multiple locations. Despite the similarities, these signals are easily identifiable as small interstellar clouds of cold hydrogen (Hi) in the galaxy. We hypothesize that the Wow! Signal was caused by a sudden brightening of the hydrogen line in these clouds triggered by a strong transient radiation source, such as a magnetar flare or a soft gamma repeater (SGR). A maser flare or superradiance mechanisms can produce stimulated emission consistent with the Wow! Signal. Our hypothesis explains all observed properties of the Wow! Signal, proposes a new source of false positives in technosignature searches, and suggests that the Wow! Signal could be the first recorded event of an astronomical maser-like flare in the hydrogen line.
Search for extraterrestrial intelligence (2127), Neutral hydrogen clouds (1099), Interstellar masers (846), Magnetars (992), Radio astronomy (1338), History of astronomy (1868)
software: The Arecibo Observatory Software Library, IDLAstro, Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Aladin Sky Atlas.
@@ 1 Introduction
The Wow! Signal, detected on August 15, 1977, by the Ohio State University’s Big Ear radio telescope, remains one of the most compelling potential instances of extraterrestrial communication (Ehman, 2011a). The signal, identified by astronomer Jerry R. Ehman, exhibited a narrowband radio frequency of approximately 1420.456 MHz, which is notably close to the hydrogen line frequency, a region of the radio spectrum often considered a natural candidate for interstellar communication (Kraus et al., 1979).
The signal lasted for 72 seconds, corresponding to the time it took the telescope beam to sweep past the source, but no follow-up observations have successfully replicated this anomaly (Gray & Ellingsen, 2002a). The transient nature of the Wow! Signal, combined with its strength and frequency, has led to extensive debate and investigation within the SETI community (Search for Extraterrestrial Intelligence), yet its exact origin remains undetermined (see, e.g., Tarter 2001).
@@
Despite numerous efforts using various radio telescopes, including the Very Large Array (VLA), no repeat detections have been made at the original frequency or any other candidate frequencies (Gray & Marvel, 2001; Backus et al., 2004; Perez et al., 2022). The search for similar signals continues to be a significant focus within the SETI community, employing advanced digital signal processing techniques and expanded observational capabilities to improve the chances of detecting faint and transient signals (Werthimer et al., 2001).
The Allen Telescope Array (ATA) was also used to search for possible repetitions (Harp et al., 2020). The researchers monitored the specific frequency and sky coordinates associated with the original signal, using the ATA’s capabilities to capture a broad frequency range in search of potential extraterrestrial signals. Despite comprehensive observation efforts spanning multiple sessions, the study did not result in the detection of any repeat signals. This outcome suggests that the Wow! Signal may have been a unique occurrence or an artifact of terrestrial interference, rather than a signal of extraterrestrial origin.
The Wow! Signal is frequently described as narrowband in contrast to the broadband spectra typical of most astronomical radiation. However, it is crucial to note that the Wow! Signal bandwidth is around 10 kHz, which is ten times the bandwidth of some astrophysical masers (Rajabi & Houde, 2020) and a thousand times greater than the bandwidth of signals targeted in numerous contemporary radio SETI searches (Wright, 2022; Huston & Wright, 2022). For example, UCLA SETI handles data with a 3 Hz resolution (Margot et al., 2023).
The main objective of this paper is to present the first results of our Arecibo Wow! project. We include a comparison between the Ohio State University (Ohio SETI) and the Arecibo Observatory SETI projects, which should provide a context of the Wow! Signal and our observations (section 2 and section 3). Arecibo Wow! is a side project of Arecibo REDS, and they are both described in section 4. We present the analysis of our observations in section 5. Our proposed mechanisms for the origin of the Wow! Signal is described in section 6. Finally, we discuss the implications of this study in section 7 and summarize our study findings in section 8.
2 Ohio SETI
@@
In 1973, following a shift in funding priorities by the United States Congress, the National Science Foundation (NSF) withdrew its financial support for the Ohio Sky Survey, which had a significant impact on the project. The loss of funding led to the cessation of the Ohio Sky Survey and the eventual disbandment of the survey team. Despite these setbacks, the Big Ear radio telescope was repurposed for a new mission: the systematic search for narrowband radio signals, or Ohio SETI.
Unlike wide-band radio sources, which are typically natural and generate emissions across a broad spectrum (including radio, optical, X-ray, and gamma-ray bands), narrowband sources are often artificial, such as those used in AM, FM, TV broadcasts, satellite communications, and radar, implying potential intelligent origin. With the narrowband search underway, a systematic survey of the declination range from -36° to +64° began (Dixon & Cole, 1977).
@@
The data collected by the Big Ear radio telescope were processed and recorded in a series of computer printouts, each row encapsulating approximately 12 seconds of sidereal time (Figure 1). During each 10-second period, the telescope collected one intensity value per second across 50 channels, averaging these ten values to represent the intensity of each channel throughout the 10-second integration period (Dixon & Cole, 1977).
The remaining 2 seconds of the 12-second interval were dedicated to computer analysis, where it processed the data and identified any notable phenomena. The printouts displayed the average intensities for each of the 50 channels (each having $\Delta\nu=10$ kHz), arranged sequentially from left to right, with channel 1 on the left side and channel 50 on the right side.
Figure 1: Full page of the computer printout with the Wow! Signal labeled in handwritten red ink. The coordinates column corresponds to the location of the positive horn. The signals are calculated from the signal-to-noise ratio (SNR) of the absolute value of the difference between the positive (ON) and negative (OFF) horn, which is 45 arcminutes ahead (3 minutes). Three locations of interest were added to the figure, labeled A, B, and C, and highlighted in red. They have the approximate azimuth size of the Big Ear telescope (8 arcminutes). The actual location of the Wow! Signal is either in B or C. If the signal was present and persistent in B, then the signal would appear in both A and B, but there is nothing in A. If present and persistent in C, then it would appear in both B and C, but there is nothing in C. A Digitized Sky Survey 2 (DSS2) frame was included for reference with the objects in the observed field (right). Printout credit: The Ohio History Connection Collections.
@@
Intensity values were converted to a single character for display: zero values were rendered as blanks, integers 1 to 9 were printed directly as digits, and integers 10 to 35 were represented by uppercase letters A through Z, respectively. Each character corresponds to a specific range of signal-to-noise ratios. During a standard analysis of the data gathered on August 15, an unusual character string, “6EQUJ5,” was detected in channel 2 of the printout on August 19, 1977.
This sequence was immediately recognized as indicative of a narrowband radio source with a small angular diameter, a potential signal of intelligent origin. The strongest signal, represented by the character “U,” indicates that the signal was 30.5 $\pm$ 0.5 times stronger than the background noise. This signal was later referred to as the Wow! Signal. In section 2 we describe its characteristics, including location, time, and absolute intensity.
@@
Frequency 1420.4556 $\pm$ 0.005 MHz
Observation Frame Galactic Center of Rest (GCR)
Bandwidth narrowband ($\leq$ 10 kHz)
Signal Strength 30.5 $\pm$ 0.5 times background noise (SNR $\approx$ 30 to 31)
@@
Positive (West) Horn Coordinates (Equatorial) RA: 19h 25m 31s $\pm$ 10s, Dec: -26d 57m $\pm$ 20m (J2000)
Negative (East) Horn Coordinates (Equatorial) RA: 19h 28m 22s $\pm$ 10s, Dec: -26d 57m $\pm$ 20m (J2000)
Positive (West) Horn Coordinates (Galactic) lon: 11d 39.0m $\pm$ 0.9m, lat: -18d 53.4m $\pm$ 2.1m
Negative (East) Horn Coordinates (Galactic) lon: 11d 54.0m $\pm$ 0.9m, lat: -19d 28.8m $\pm$ 2.1m
@@
The hypothesis that has gained the most attention, especially among the public, is that the Wow! Signal was a transmission from an extraterrestrial civilization (Wright, 2021). The frequency of the signal, close to the hydrogen line, is considered a logical choice for interstellar communication because it is naturally quiet and would be universally recognized by technologically advanced civilizations (Drake & Sagan, 1973). However, this hypothesis faces critical issues.
The signal was never repeated, despite numerous follow-up observations. In the context of SETI, repetition is essential for verification, as a one-time detection makes it difficult to rule out other sources. Moreover, the absence of modulation or encoding in the signal, which would be expected from an intentional transmission, further complicates this interpretation.
@@
Another possibility is that the Wow! Signal originated from a natural astrophysical source. Candidates include stellar flares, pulsars, or other cosmic phenomena that emit radio waves (for a recent explanation overview, see Kipping & Gray 2022). However, these sources typically exhibit characteristics such as periodicity or broader bandwidths, which the Wow! Signal lacks.
The narrowband nature of the signal is particularly unusual because natural emissions are generally broadband. Additionally, the hydrogen line frequency is not usually associated with strong (non-thermal) radio emissions from astrophysical objects, making this hypothesis less likely.
The possibility of interstellar scintillation (ISS) has also been raised as a potential explanation for the Wow! Signal. In the case of ISS, a radio signal can fluctuate significantly in strength and phase when passing through the ionized interstellar medium (Cordes & Lazio, 1991; Cordes et al., 1997; Brzycki et al., 2023). A key obstacle is that ISS will not make a broadband signal narrowband and the probability of a scintillation gain on the order of a hundred is very small (Gray & Marvel, 2001).
@@
Terrestrial interference is another plausible explanation. The signal could have been caused by a reflected terrestrial transmission or an Earth-based signal that temporarily overwhelmed the receiver. The 1420 MHz band is protected for astronomical purposes, but illegal or accidental transmission in this band could have occurred. However, the Big Ear radio telescope had a design that helped filter out terrestrial sources, and the signal appeared only in one of the two feed horns, complicating the explanation of interference. No known terrestrial source has been identified that matches the characteristics of the Wow! Signal. [^1] For a description and some of the hypotheses that have been raised about the terrestrial and extraterrestrial origin of the signal see the detailed account written by Jerry R. Ehman himself in 2010 at this URL (last visited: September 11, 2024).
@@
One hypothesis suggests that the signal could have been caused by a reflection of Earth-based signals off space debris or satellites (Ehman, 2011b). This idea posits that a radio signal from Earth might have bounced off a satellite or space debris and was then picked up by the telescope. However, this explanation struggles with the same issue of non-repetition.
The duration of the signal and its specific characteristics do not align neatly with the known satellite reflections or the behavior of the debris. Additionally, the Doppler shift expected from such a reflection was not observed. Moreover, a reflecting piece of space debris must have been moving very slowly and not tumbling to mimic a point source in the celestial sphere, which is unlikely (Ehman, 2011a).
@@
There is also the consideration of interstellar propagation effects, where the signal could have been distorted or focused through gravitational lensing or other space phenomena. These effects could potentially create a signal that mimics the characteristics of the Wow! Signal. However, the gravitational lensing hypothesis suffers from the drawback of a short signal duration. The Wow! Signal was not detected in the second horn of the Big Ear whose beam swept the same area of the sky only 3 minutes later (Ehman, 2011a). Additionally, such a scenario would likely involve complex interactions with multiple cosmic objects, which are improbable and would likely have been observed again under similar conditions.
@@
Finally, human error has been considered in the recording or processing of the data. Data handling methods in 1977 were not as advanced as today, and it is conceivable that a mistake or anomaly in the data processing chain produced the Wow! Signal. However, this explanation lacks specific evidence and does not account for the signal’s characteristics, which seem consistent with a real astronomical event.
In summary, any explanation of the Wow! Signal must account for its five principal characteristics (Ehman, 2010): (1) it exhibited a narrowband nature, with a bandwidth equal or less than 10 kHz, (2) the mean signal strength remained invariant for a minimum of 72 seconds, persisting for a duration ranging from minutes to hours, as it was detected in only one of two sequential horns (separated by a three-minute interval) and not within a 24-hour period succeeding this event, (3) it possessed a high flux density, measured in the range of tens to hundreds of Janskys, (4) no modulation was discerned on a temporal scale shorter than 10 seconds, and (5) it was a singular observation, not observed in subsequent observations by Big Ear or any other telescopes henceforth.
@@
Table 2: Comparison of the telescopes of the Arecibo Observatory (AO) and the Ohio State University Radio Observatory (OSURO).
@@
Telescope Type Fixed dish with moving receiver Fixed paraboloid with meridian-transit
Operational Years 1963 - 2020 1963 - 1998
Location Arecibo, Puerto Rico, USA Delaware, Ohio, USA
@@
Elevation 497 m 276 m
Main Reflector Spherical Reflector Cylindrical Paraboloid
Dimensions 305 m diameter 104 m wide by 21 m height
Equivalent Dish Diameter 305 m 53 m
Frequency Range 300 MHz to 10 GHz 1411 to 1419 MHz (also 612 and 2650 MHz)
Wavelength Rage 1 m to 3 cm 21 cm (also 49 cm and 11 cm)
Illuminated Area 213 m x 237 m (40,000 m2) $\sim$2,200 m2
Effective Dish Diameter 225 m $\sim$53 m
Radius of Curvature of Primary Reflector 265 m $\sim$243 m
Maximum Declination Coverage -1 to +38 degrees -36 to +63 degrees
@@
Zenith Angle Range 40 degrees 100 degrees
Pointing Accuracy $\sim$1 arcsecond unknown
References. — (Ehman, 2011a; Salter, 2012)
Table 3: Comparison of the observation parameters of Arecibo Wow!, relevant to this study, with those of the Ohio SETI.
Property Arecibo Wow! Ohio SETI
Operation Years May 2024 - Present 1973 to 1995 (22 years)
Telescope Arecibo Telescope OSU Telescope (Big Ear)
Band L-band L-band
@@
Frequency Range 1164.5 to 1735.5 MHz 1420.4 to 1420.9 MHz
Frequency Channels 7 x 8192 50aaOhio SETI operated with 8 channels, 20 to 100 kHz bandwidth, from 1973 to 1976.
Bandwidth/channel 10.1 kHz 10 kHz
Total Bandwidth 571 MHz 500 kHz
Integration Time 0.1 seconds 12 secondsbbThe integration time was 10 seconds, plus a 2-second wait time between integrations.
Polarization Full Stokes Single
Observation Modes ON, ON-OFF, DRIFT DRIFT
Backend Mock Spectrometer Ohio SETI
Gain 7.9 K/JyccPre-Hurricane Maria values were $\sim$10.5 K/Jy and $\sim$2.4 Jy, respectively. $\sim$0.1 K/Jy
System Equivalent Flux Density (SEFD) 3.2 JyccPre-Hurricane Maria values were $\sim$10.5 K/Jy and $\sim$2.4 Jy, respectively. $\sim$1000 Jy
Half-Power Beam Width (HPBW) (Az x ZA) 3.1 x 3.5 arcminutes 8 x 40 arcminutes
Transit Time (null-to-null) @ ($\pm$27° declination) 28 seconds 72 seconds
Note. — Az = azimuth, ZA = zenith angle.
References. — (Kraus et al., 1979; Dixon, 1985; Salter, 2012)
3 Arecibo SETI